Circular No. 7336

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gn IN M61 K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka,

Kyushu University, report that a wide-band spectrum (range 400–800 nm; resolution 0.6 nm) of SN 1999gn, obtained with the BAO 1.01-m reflector on Dec. 20.8 UT, shows it to be a type-II supernova in an early phase. It exhibits a blue continuum with a broad P-Cyg Hα profile and Hβ absorption. The expansion velocity of the supernova, derived from Hα absorption minimum with respect to the narrow Hα emission (from the superposed H II region), is 5300 km/s.

SUPERNOVA 1999em IN NGC 1637 C. K. Lacey, Naval Research Laboratory (NRL); S. D. Van Dyk, IPAC,

California Institute of Technology; K. W. Weiler, NRL; R. A. Sramek, National Radio Astronomy Observatory (NRAO); and N. Panagia, Space Telescope Science Institute, report: “Radio observations with the NRAO Very Large Array on Dec. 1 yield the detection of a compact radio source at 8.4 GHz (3.6 cm) of 0.19 ± 0.03 mJy at position α = 4^{h}41^{m}27^{s}.157, δ = −2^{o}51′45′′.83 (equinox 2000.0, uncertainty ± 0′′.1) displaced ∼ 1′′.5 in α from the precise optical position reported on IAUC 7296. However, a new measurement of the optical position by Van Dyk from an R-band image of SN 1999em made with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Nov. 5 UT, and kindly provided by W. D. Li and A. V. Filippenko, yields optical position end figures 27^{s}.13, 45′′.4, which agrees to within 0′′.6 with the radio position. Thus, this appears to represent a detection of radio emission from the type-IIP supernova 1999em with an 8.4-GHz spectral luminosity of 1.4 × 10^{25} erg s^{−1} Hz^{−1} for an assumed distance of 7.8 Mpc (Sohn and Davidge 1998, A.J. 115, 130). On Dec. 13, a new measurement at 8.4 GHz yields a flux density of 0.22 ± 0.04 mJy. The full set of radio measurements, including upper limits reported earlier on IAUC

7318, is available at http://rsd-www.nrl.navy.mil/7214/weiler/sne-
home.html.”

GRS 1915+105 E. B. Waltman, Naval Research Laboratory; and F. D. Ghigo, National

Radio Astronomy Observatory, report: “The Green Bank Interferometer has observed a flare in the microquasar GRS 1915+105 on Dec. 23. Flux densities reached 0.5 Jy at 2 GHz on Dec. 23.8 UT, ending the current radio plateau state. The data may be accessed at http://www.gb.nrao.edu.”

1999 December 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green